Q(X,y):X+y >or=7 then what is truth value of∀x∃yQ(x,y)


1
Expert's answer
2021-02-22T13:11:41-0500

 The answer depends on the set where the values of x, y are defined. Given ,

"Q(x,y):x+y\ge7"


"\forall x \exist yQ(x, y)"


For all value of x Let x=1 and y=1:



"x+y=1+1=2<7""2\text{ is less than }7"

Since "x+y\le7" The truth value of Q is false for some values of x and y.


"\text{At } x=8,y=2\\

X+y=8+2=10>7"


then the truth value of Q(x, y) is true for some values of x and y